A vertical storage tank with a hemispherical bottom and
a cylindrical shell of 2 m. inside diameter and 4 m. high
is filled with water. A sharp-edge orifice 150 mm
diameter with C = 0.62 is located at the lowest point.
26. Determine the time to empty the cylindrical part of
the tank through the orifice.
A. 134.04 sec C. 152.11 sec
B. 145.6 sec D. 160 sec
27. Determine the time to empty the hemispherical
tank only through the orifice.
A. 60.41 sec C. 75.11 sec
B. 53.61 sec D. 48.11 sec
28. Determine the time to empty the whole tank
through the orifice.
A. 213.61 sec C. 220.41 sec
B. 187.65 sec D. 194.45 sec
Answer #1

Solution - Giren - R> zim. * 2m=D dia oforifice Cat 150mm = 0.15 m. (a=0.62. 4m im orifice of dia = 150mm o orifice area =a=1 dt -TRY dz (t-o) = -TIR Qidt = -1Ř. dzi relo cirty through orifice J2qz -:Q = Ca a vegz where a = area of sufice (da. Vagz.w.kit R=1m; (a = 0.62 a= Tx0.15 mr. g=9.81 m/sea Z,= 5m! Zz=im. it= 2ft xiv 0.62x + x 0.15 x/2x9.81 4 => t = 8. 0-62X0-15 x 1we know that do Q.at - dv here. – dv=-A.dz - Tar, dz. = T[R_(R-Z] dz =- [R² _ [2 +2% -2R2)] dz =-(-_Z +2Rz) az Far= T[=_282)For this case ZER=in 2. = om 2252 L-03- GR. Z 3 ( a. Izg t- -|(2022 48.64)- Ca a 129 ( 2XR31_UR-p? 3 T Ga Jeg [04 aeth 4 esh)